Can anyone answer this?
Fifty students were asked their favorite season.Thirty saaid summer,15 said winter and the rest spring or fall.In a circle graph what would the measure of each central angle be?

Answer by rapaljer(4551)

Summer = 30 out of 50 which is 60%
Winter = 15 out of 50 which is 30%

The rest would be spring or fall, which is 50 - 45 = 5.
Spring or Fall = 5 out of 50 which is 10%.

In a circle, the entire circle represents 360 degrees.
Summer = 60% of 360 = .6*360 = 216 degrees
Winter = 30% of 360 = .3*360 = 108 degrees
Spring or fall = 10% =.1*360 = 36 degrees
Check: Add these together and see if they add up to 360 degrees. They do, so it checks.

R^2 at SCC